OpenStudy (anonymous):

If f(x)=x^2 + 2, find and simplify the following... f(t+1)=______ ^^ would it be like this? f(t+1)=(t+1)^2 + 2 ? if so, would it need to be simplified more? f(t^2 + 1) = ______ ^^would it be like this? f(t^2 + 1)=((t^2 + 1)^2+1) again, would it need to be simplified more? f(2)=_____ ^^would it be f(2)=2^2 + 2 = 4+2=6 ?? 2f(t)=____ ^^not too sure on this one.... would it start off like this? 2(t^2+2)=2t^2 + 4 ?? is that right? :/ (f(t))^2 + 1 =_______ ^^not very sure on this one.. :( Please explain? Thanks!!! :)

OpenStudy (agent0smith):

^yeah in that case, usually simplifying is better.

OpenStudy (anonymous):

ohh okay... so for the first one it just becomes t^2 + 1 +2 = t^2 + 3 ?

OpenStudy (agent0smith):

You overcomplicated this one: 2f(t). Find f(t) first, that's easy. Then multiply it by 2.

OpenStudy (anonymous):

ohh okay... so 2f(t)=(t^2+2)2 = 2t^2 + 4 ?? is that right?

OpenStudy (agent0smith):

Oh you kinda did that, just your explanation was weird. Yes.

OpenStudy (anonymous):

haha ohh oops sorry about that :P

OpenStudy (agent0smith):

(f(t))^2 + 1 Square f(t), and then add 1

OpenStudy (anonymous):

(t^2+2)^2= t^4 + 4 ? then add 1, and would you get t^4 + 5 ? is that correct? :/

OpenStudy (agent0smith):

^nope. You can not do that. (a+b)^2 does NOT equal a^2 + b^2 How do you square a number? Multiply it by itself. Do that with (t^2+2)^2

OpenStudy (anonymous):

aw darn... so 2t^2 + 4 ? :/

OpenStudy (agent0smith):

a^2 = a*a (a+b)^2 = (a+b)(a+b)

OpenStudy (anonymous):

ohh i see now so (t^2+2)(t^2+2)= t^4 + 4t^2 + 4 ??

OpenStudy (agent0smith):

Now find (f(t))^2 + 1

OpenStudy (anonymous):

t^4 + 4t^2 + 5 ?

OpenStudy (anonymous):

if so, are these the correct answers? f(t+1)= t^2 + 2t + 4 f(t^2+1) = t^4 + 2t^2 + 3 f(2) = 6 2f(t) = 2t^2 + 4 (f(t))^2 + 1 = t^2 + 2t + 4 are those all correct? i fixed the first two and did it the way you said using FOIL... :/

OpenStudy (agent0smith):

(f(t))^2 + 1 = t^2 + 2t + 4 you have this wrong now, had the right answer above.

OpenStudy (anonymous):

oh oops sorry i meant + 5 haha (f(t))^2 + 1 = t^2 + 2t + 5 so when it's 5, are the all correct now? :)

OpenStudy (agent0smith):

Probably, yes.

OpenStudy (anonymous):

okay awesome!! :) yay :) thanks so much!!! :D